OpenStudy (anonymous):

Let u = <-4, -3>. Find the unit vector in the direction of u, and write your answer in component form Help me understand how to do this problem?

OpenStudy (anonymous):

?

OpenStudy (anonymous):

Looks like you and @gloves are in the same class?

OpenStudy (anonymous):

He have a similar question?

OpenStudy (anonymous):

We want to find \(c\) such that we multiply it by \(\mathbf u\) and have a vector magnitude 1. In short: \(|c\mathbf u|=1\).\[ |c\mathbf u| = \sqrt{c\mathbf u \cdot c\mathbf u} = \sqrt{c^2(\mathbf u \cdot \mathbf u)} = c|\mathbf u| \implies c = \frac{1}{|\mathbf u|} \]Therefore, the unit vector for \(\mathbf u\) is \(\mathbf u / |\mathbf u|\)

OpenStudy (anonymous):

The reason why I did that, is because multiplying a vector by a scalar preserves direction, but not magnitude.

OpenStudy (anonymous):

The unit vector has the same direction, but magnitude 1.

OpenStudy (anonymous):

Ok so what number do I plug into that?

OpenStudy (anonymous):

Well, \(\mathbf u = \langle -4,-3\rangle\)

OpenStudy (anonymous):

Yeah so both -4 and -3?

OpenStudy (anonymous):

These are the answers

OpenStudy (anonymous):

First, do you know the dot product? \(\mathbf u \cdot \mathbf u\)?

OpenStudy (anonymous):

No

OpenStudy (anonymous):

Wait, really?

OpenStudy (anonymous):

Wasn't in the lesson, the next lesson includes it

OpenStudy (anonymous):

Okay. Do you know how to get the angle for \(\mathbf u\)?

OpenStudy (anonymous):

No :(

OpenStudy (anonymous):



OpenStudy (anonymous):

wait i got it

OpenStudy (anonymous):



OpenStudy (anonymous):

I looked at this example

OpenStudy (anonymous):

Yes, but you can't really use that if you don't know the dot product yet.

OpenStudy (anonymous):

Wait, did you learn \(|\mathbf u|\)? That is how to find the magnitude of a vector?

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

Ok.... Then you should have understood when I said \[ \frac{\mathbf u}{|\mathbf u|} \]